Bill invests in a plan that has an interest rate of 5%. He invests two times as much in a plan that has an interest rate of 6%. If the total interest from the investments is $357 after one year, how much was invested in each plan?
Answer by addingup(3677)

0.05x+0.06(2x) = 357
0.05x+0.12x = 357
0.17x = 357
x = 2,100 this is the amount invested at 5%
The amount at 6% is 2*2,100 = 4,200